# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v3

@@ VI.6 PYATB
PYATB (PYthon Ab initio Tight-Binding simulation package) is a Python package based on ab initio tight-binding Hamiltonian, designed as a tool for calculating and analyzing the electronic structures of materials Jin et al. (2023). It can be viewed as a post-processing program for ABACUS. When ABACUS completes the SCF calculations and generates the tight-binding Hamiltonian, PYATB utilizes this Hamiltonian to perform electronic property calculations. This eliminates the need for the cumbersome construction of MLWFs while strictly preserving the Hamiltonian’s symmetry, making it particularly well-suited for high-throughput workflows in electronic structure calculations of various materials. Currently, PYATB mainly offers three functional modules, i.e., Bands, Geometric, and Optical. This allows ABACUS, in combination with PYATB, to compute basic band structures, fat bands, and projected density of states (PDOS). Additionally, it can analyze the Berry curvature and Chern number of topological materialsXiao, Chang, and Niu (2010); Hasan and Kane (2010), as well as compute first- and second-order optical responses such as optical conductivity, shift current, second harmonic generation Sturman and Fridkin (1992); Sipe and Shkrebtii (2000); Jin and He and Berry curvature dipoleSodemann and Fu (2015); Pang, Jin, and He .
@@
VI.7 Hefei-NAMD
Hefei Non-Adiabatic Molecular Dynamics (Hefei-NAMD) is an ab initio simulation suite for studying excited carrier dynamics in condensed matter systems. It combines real-time time-dependent density functional theory (TDDFT) with the fewest-switches surface hopping scheme and classical-path approximation. Hefei-NAMD has been used to investigate processes such as charge transfer, electron–hole recombination, spin dynamics, and exciton dynamics Zheng et al. (2019); Zheng, Zheng, and Zhao (2022); Jiang et al. (2021); Zheng et al. (2023). It allows the study of excited carrier dynamics in energy, real, and momentum spaces, while also exploring interactions with phonons, defects, and molecular adsorptions, providing valuable insights into ultrafast dynamics at the atomic scale. Hefei-NAMD works by interfacing with other ab initio codes, e.g. VASP, ABACUS, etc. Figure 34 shows the flowchart of the simulation using Hefei-NAMD interfaced with ABACUS.
@@ VI.8 PEXSI
Notably, PEXSI exhibits superior computational scaling compared to traditional diagonalization methods. While standard diagonalization scales cubically as $O(N^{3})$, PEXSI achieves at most quadratic scaling ($O(N^{2})$), where $N$ is the dimension of the matrix The method’s architecture features a two-level parallelization scheme, incorporating both pole expansion parallelism and selected inversion parallelism, thereby demonstrating exceptional scalability. ABACUS has implemented an interface to the parallel version of PEXSI for solving the KS equation expanded with the NAO basis.
VI.9 Other Packages
ASE (Atomic Simulation Environment) comprises a comprehensive suite of tools and Python modules designed for setting up, manipulating, executing, visualizing, and analyzing atomic simulations ASE (2024a). We have developed an ABACUS calculator ASE (2024b) that interfaces with ASE (version 3.23.0b1) and that enables seamless integration of ASE’s robust functionalities for both pre-processing and post-processing tasks. Regarding pre-processing capabilities, the interface facilitates tasks such as converting structural files into various formats, generating $k$-point grids, etc. For post-processing, the integration with ASE empowers ABACUS users with advanced tools for visualizing and analyzing simulation results. Moreover, the self-consistent calculations of ABACUS can be combined with ASE’s built-in optimization algorithms, facilitating structural relaxation procedures, conducting MD simulations under varying conditions, performing precise phonon calculations based on the finite displacement method, carrying-out global optimization by genetic algorithm (GA), locating transition states (TS) of reactions by saddle point refinement, etc. For TS search, as an example, the nudged elastic band (NEB) Jónsson, Mills, and Jacobsen (1998); Henkelman and Jónsson (2000); Henkelman, Uberuaga, and Jónsson (2000) method and the Dimer Henkelman and Jónsson (1999); Olsen et al. (2004); Heyden, Bell, and Keil (2005); Kästner and Sherwood (2008) method, which represent double-ended TS searching methods and single-end TS methods respectively, are implemented in ASE. Besides, some other enhanced NEB methods are coded in the ASE package, such as dynamic NEB Lindgren, Kastlunger, and Peterson (2019) and AutoNEB Kolsbjerg, Groves, and Hammer (2016). Furthermore, there are other saddle point refinement algorithms implemented in ASE, and some of them show better performance, like Sella Hermes et al. (2019, 2021, 2022). With the above-mentioned methods available in ASE, ABACUS can be used for searching for transition states.
@@ VI.9 Other Packages
TB2JHe et al. (2021) is an open-source Python package for automatic computation of magnetic interactions between atoms of magnetic crystals from density functional Hamiltonians based on Wannier functions or linear combinations of atomic orbitals. The program is based on Green’s function method with the local rigid spin rotation treated as a perturbation Liechtenstein et al. (1987). The ABACUS interface has been added since TB2J version 0.8.0. With the Hamiltonian matrix $\mathbf{H}(\mathbf{k})$ and the overlap matrix $\mathbf{S}(\mathbf{k})$ obtained by an LCAO calculation, TB2J is able to compute the isotropic exchange, symmetric anisotropic exchange, and the Dzyaloshinskii-Moriya interaction (DMI). In 2024, Zhang et al. Zhang et al. (2024e) study GeFe3N by using ABACUS-TB2J, which help them study the short-range order and strong interplay between local and itinerant magnetism in GeFe3N.
Wannier90 Pizzi et al. (2020) is a software designed for calculating maximally-localized Wannier functions (MLWFs) Marzari and Vanderbilt (1997), which facilitates cost-effective band structure calculations. MLWFs are often used to construct model Hamiltonians. The post-processing capabilities of Wannier90 enable the calculation of various properties, such as Berry curvature, Berry curvature dipole, and shift currentPizzi et al. (2020). Additionally, MLWFs generated by Wannier90 can be analyzed for material surface states and topological properties using the Wannier Tools Wu et al. (2018). According to the standard scheme for generating maximally localized Wannier functions in Wannier90, ABACUS has implemented an interface compatible with the Wannier90 software. This interface allows ABACUS to generate the necessary files for Wannier90, including overlap files between the periodic parts of Bloch functions at neighboring $\mathbf{k}$-points, projection files of Bloch functions onto trial localized orbitals, eigenvalue files of the energy bands, and real-space distribution files of the periodic parts of Bloch functions.
@@ VII Summary
ABACUS has been used to generate a large amount of first-principles data for training machine-learning-based potentials. The APNS project has been initiated to ensure the reliability and accuracy of ABACUS, where systematic tests for pseudopotentials and numerical atomic orbitals have been conducted, providing a solid foundation for the software’s performance. ABACUS interacts with several packages, such as DeePKS-kit, DeePMD-kit, DP-GEN, DeepH, DeePTB, PyATB, Hefei-NAMD, PEXSI, etc., further expanding its functionality and applicability in multi-scale calculations and electronic structure analysis.
@@ Acknowledgements.
We thank Han Wang and Weinan E for many helpful discussions. M.C. gratefully acknowledges AI for Science Institute, Beijing (AISI), as well as funding support from the National Key R&D Program of China under Grant No.2025YFB3003603 and the National Natural Science Foundation of China (grant no. 11988102, 12135002, 12122401,12074007). S.X. gratefully acknowledges funding support from the National Natural Science Foundation of China (grant no. 52273223), DP Technology Corporation (grant no. 2021110016001141), School of Materials Science and Engineering at Peking University, and the AI for Science Institute, Beijing (AISI). S.L. acknowledges the supports from National Natural Science Foundation of China (92370104) and Natural Science Foundation of Zhejiang Province (2022XHSJJ006). The computational resource is provided by Westlake HPC Center. M.L. gratefully acknowledges the financial support of the National Natural Science Foundation of China (No. 52276212), National Key Research and Development Program of China (No. 2022YFB3803600). Q.O. acknowledges research grants from China Petroleum &\& Chemical Corp (funding number 124014). T.W. acknowledges support by The University of Hong Kong (HKU) via seed funds (2201100392, 2409100597). Y.X. acknowledges support by the National Key Basic Research and Development Program of China (grant nos. 2024YFA1409100 and 2023YFA1406400), the National Natural Science Foundation of China (grants nos. 12334003, 12421004, and 12361141826), and the National Science Fund for Distinguished Young Scholars (grant no. 12025405). W.Z. gratefully acknowledges AI for Science Institute, Beijing (AISI), as well as support from the Hongyi postdoctoral fellowship of Wuhan University.

### v4

@@ VI.7 PYATB
PYATB (PYthon Ab initio Tight-Binding simulation package) is a Python package based on ab initio tight-binding Hamiltonian, designed as a tool for calculating and analyzing the electronic structures of materials. Jin et al. (2023) It can be viewed as a post-processing program for ABACUS. When ABACUS completes the SCF calculations and generates the tight-binding Hamiltonian, PYATB utilizes this Hamiltonian to perform electronic property calculations. This eliminates the need for the cumbersome construction of Maximally localized Wannier functions (MLWFs) while strictly preserving the Hamiltonian’s symmetry, making it particularly well-suited for high-throughput workflows in electronic structure calculations of various materials. Currently, PYATB mainly offers three functional modules, i.e., Bands, Geometric, and Optical. This allows ABACUS, in combination with PYATB, to compute basic band structures, fat bands, and projected density of states (PDOS). Additionally, it can analyze the Berry curvature and Chern number of topological materials, Xiao, Chang, and Niu (2010); Hasan and Kane (2010) as well as compute first- and second-order optical responses such as optical conductivity, shift current, second harmonic generation Sturman and Fridkin (1992); Sipe and Shkrebtii (2000); Jin and He and Berry curvature dipole. Sodemann and Fu (2015); Pang, Jin, and He
@@
VI.8 Hefei-NAMD
Hefei Non-Adiabatic Molecular Dynamics (Hefei-NAMD) is an ab initio simulation suite for studying excited carrier dynamics in condensed matter systems. It combines real-time time-dependent density functional theory (TDDFT) with the fewest-switches surface hopping scheme and classical-path approximation. Hefei-NAMD has been used to investigate processes such as charge transfer, electron–hole recombination, spin dynamics, and exciton dynamics. Zheng et al. (2019); Zheng, Zheng, and Zhao (2022); Jiang et al. (2021); Zheng et al. (2023) It allows the study of excited carrier dynamics in energy, real, and momentum spaces, while also exploring interactions with phonons, defects, and molecular adsorptions, providing valuable insights into ultrafast dynamics at the atomic scale. Hefei-NAMD works by interfacing with other ab initio codes, e.g. VASP, ABACUS, etc. Fig. 33 shows the flowchart of the simulation using Hefei-NAMD interfaced with ABACUS.
@@ VI.9 PEXSI
Notably, PEXSI exhibits superior computational scaling compared to traditional diagonalization methods. While standard diagonalization scales cubically as $O(N^{3})$, PEXSI achieves at most quadratic scaling ($O(N^{2})$), where $N$ is the dimension of the matrix. The method’s architecture features a two-level parallelization scheme, incorporating both pole expansion parallelism and selected inversion parallelism, thereby demonstrating exceptional scalability. ABACUS has implemented an interface to the parallel version of PEXSI for solving the KS equation expanded with the NAO basis.
VI.10 Other Packages
ASE (Atomic Simulation Environment) comprises a comprehensive suite of tools and Python modules designed for setting up, manipulating, executing, visualizing, and analyzing atomic simulations. ASE (2024a) We have developed an ABACUS calculator ASE (2024b) that interfaces with ASE (version 3.23.0b1) and that enables seamless integration of ASE’s robust functionalities for both pre-processing and post-processing tasks. Regarding pre-processing capabilities, the interface facilitates tasks such as converting structural files into various formats, generating $k$-point grids, etc. For post-processing, the integration with ASE empowers ABACUS users with advanced tools for visualizing and analyzing simulation results. Moreover, the self-consistent calculations of ABACUS can be combined with ASE’s built-in optimization algorithms, facilitating structural relaxation procedures, conducting MD simulations under varying conditions, performing precise phonon calculations based on the finite displacement method, carrying-out global optimization by genetic algorithm (GA), locating transition states (TS) of reactions by saddle point refinement, etc. For TS search, as an example, the nudged elastic band (NEB) Jónsson, Mills, and Jacobsen (1998); Henkelman and Jónsson (2000); Henkelman, Uberuaga, and Jónsson (2000) method and the Dimer Henkelman and Jónsson (1999); Olsen et al. (2004); Heyden, Bell, and Keil (2005); Kästner and Sherwood (2008) method, which represent double-ended TS searching methods and single-end TS methods respectively, are implemented in ASE. Besides, some other enhanced NEB methods are coded in the ASE package, such as dynamic NEB Lindgren, Kastlunger, and Peterson (2019) and AutoNEB. Kolsbjerg, Groves, and Hammer (2016) Furthermore, there are other saddle point refinement algorithms implemented in ASE, and some of them show better performance, like Sella. Hermes et al. (2019, 2021, 2022) With the above-mentioned methods available in ASE, ABACUS can be used for searching for transition states.
@@ VI.10 Other Packages
TB2JHe et al. (2021) is an open-source Python package for automatic computation of magnetic interactions between atoms of magnetic crystals from density functional Hamiltonians based on Wannier functions or linear combinations of atomic orbitals. The program is based on Green’s function method with the local rigid spin rotation treated as a perturbation. Liechtenstein et al. (1987) The ABACUS interface has been added since TB2J version 0.8.0. With the Hamiltonian matrix $\mathbf{H}(\mathbf{k})$ and the overlap matrix $\mathbf{S}(\mathbf{k})$ obtained by an LCAO calculation, TB2J is able to compute the isotropic exchange, symmetric anisotropic exchange, and the Dzyaloshinskii-Moriya interaction (DMI). In 2024, Zhang et al. Zhang et al. (2024f) study GeFe3N by using ABACUS-TB2J, which help them study the short-range order and strong interplay between local and itinerant magnetism in GeFe3N.
Wannier90 Pizzi et al. (2020) is a software designed for calculating maximally-localized Wannier functions (MLWFs), Marzari and Vanderbilt (1997) which facilitates cost-effective band structure calculations. MLWFs are often used to construct model Hamiltonians. The post-processing capabilities of Wannier90 enable the calculation of various properties, such as Berry curvature, Berry curvature dipole, and shift current. Pizzi et al. (2020) Additionally, MLWFs generated by Wannier90 can be analyzed for material surface states and topological properties using the Wannier Tools. Wu et al. (2018) According to the standard scheme for generating maximally localized Wannier functions in Wannier90, ABACUS has implemented an interface compatible with the Wannier90 software. This interface allows ABACUS to generate the necessary files for Wannier90, including overlap files between the periodic parts of Bloch functions at neighboring $\mathbf{k}$-points, projection files of Bloch functions onto trial localized orbitals, eigenvalue files of the energy bands, and real-space distribution files of the periodic parts of Bloch functions.
@@ VII Summary
ABACUS has been used to generate a large amount of first-principles data for training machine-learning-based potentials. The APNS project has been initiated to ensure the reliability and accuracy of ABACUS, where systematic tests for pseudopotentials and numerical atomic orbitals have been conducted, providing a solid foundation for the software’s performance. ABACUS interacts with several packages, such as DeePKS-kit, DeePMD-kit, DP-GEN, DeepH, DeePTB, HamGNN, PyATB, Hefei-NAMD, PEXSI, etc., further expanding its functionality and applicability in multi-scale calculations and electronic structure analysis.
@@ Acknowledgements.
We thank Han Wang and Weinan E for many helpful discussions. We gratefully acknowledges supports from AI for Science Institute, Beijing (AISI). M.C. thank funding support from the National Key R&D Program of China under Grant No.2025YFB3003603, the National Natural Science Foundation of China (No. 12588201, 12588301, 12135002, 12122401,12074007). H.R. is supported by Beijing Natural Science Foundation (No. QY24014). S.X. gratefully acknowledges funding support from the National Natural Science Foundation of China (grant no. 52273223), DP Technology Corporation (grant no. 2021110016001141), School of Materials Science and Engineering at Peking University, and the AI for Science Institute, Beijing (AISI). S.L. acknowledges the supports from National Natural Science Foundation of China (92370104) and Natural Science Foundation of Zhejiang Province (2022XHSJJ006). The computational resource is provided by Westlake HPC Center. M.L. gratefully acknowledges the financial support of the National Natural Science Foundation of China (No. 52276212), National Key Research and Development Program of China (No. 2022YFB3803600). Q.O. acknowledges research grants from China Petroleum &\& Chemical Corp (funding number 124014). T.W. acknowledges support by The University of Hong Kong (HKU) via seed funds (2201100392, 2409100597). Y.X. acknowledges support by the National Key Basic Research and Development Program of China (grant nos. 2024YFA1409100 and 2023YFA1406400), the National Natural Science Foundation of China (grants nos. 12334003, 12421004, and 12361141826), and the National Science Fund for Distinguished Young Scholars (grant no. 12025405). W.Z. gratefully acknowledges AI for Science Institute, Beijing (AISI), as well as support from the Hongyi postdoctoral fellowship of Wuhan University.